OpenStudy (anonymous):

In bisection search example (lecture 3), Guttag asks the class 'how large was the search space?' The answer was 'roughly x / epsilon **2'. How does one arrive at this answer?

OpenStudy (e.mccormick):

Well, how many items get searched and why is it that number? That defines the search space.

OpenStudy (anonymous):

I get why the search space would be at least x / epsilon, and that it ought to be resolved even more finely than that. Is x / epsilon**2 the 'real' answer? Or would it be as correct to have chosen something like x / (2*epsilon)? The code is: x = 12345 epsilon = .01 numGuesses = 0 low = 0.0 high = x ans = (high + low)/2.0 while (ans**2 - x) >= epsilon and ans <= x: #print low, high, ans numGuesses += 1 if ans**2 < x: low = ans else: high = ans ans = (high + low)/2.0 print ‘numGuesses = ‘, numGuesses print ans, ‘is close to square root of’, x

OpenStudy (anonymous):

Sorry, just ignore the print statements and question marks, they aren't important. I must have pasted something with the wrong encoding.

OpenStudy (e.mccormick):

Each time you search half as much, so your search soace is expotentially shrinking.